sin x 15. Simplify the following expression: (arctan Vr - 1). (Assume that x>1. of the function - 1
Answer #1

15. see Carc tan V221) Lel, arcfton V x2-1) = V x2-1 = tan c tan a = 1221 Vivae. )2 + 1 = x Vx2_1 See (arctam 1921) = see o s